OpenStudy (anonymous):

For l(t)=(5sint, 5cost) describing the motion of a circle x^2+y^2=25, what would the velocity be at t=pi/3

OpenStudy (jamesj):

Calculate the vector velocity: note that l(t) = (x(t),y(t)), then vector velocity dl/dt = (dx/dt, dy/dt)

OpenStudy (anonymous):

= magnitude of (5cos(t), -5sin(t))

OpenStudy (anonymous):

= sqrt (25)= 5m/s

OpenStudy (jamesj):

The questions asks for velocity, not speed.

OpenStudy (anonymous):

5/2,-5sqrt(3)/2

OpenStudy (jamesj):

Hm, velocity at t, v(t) = 5(cos t, -sin t). So v(pi/3) = 5( cos(pi/3), -sin(pi/3)) = ...

OpenStudy (jamesj):

Yes, you've got it.